### v1

Abstract
We find that the spontaneous and collective emissions have a large influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle. The spontaneous and collective emissions restrict the absorption enhancement by plasmonic mode and make the enhancement possible only with a fast, on a pico-second time scale, population relaxation of the upper absorbing states. Conditions for the maximum of the plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found. We show that the nanoparticle-absorber interaction and the collective emission cause the bistability in the plasmon-enhanced absorption.
I Introduction
The localized plasmon resonances (LPR) excitation leads to the electric field enhancement near metal nanoparticles Maier (2007); Pelton, Aizpurua, and Bryant (2008); Amendola et al. (2017); Tanabe (2008). Such a "plasmonic" field enhancement finds various applications in optical sensors Li, Cushing, and Wu (2015), bio-medicine Liao, Nehl, and Hafner (2006), photovoltaics Jang et al. (2016), photo-catalysis Zhang et al. (2013) and other areas of physics and chemistry Maier (2007); Barbillon (2019). A new branch of chemistry - plasmonic chemistry was developed concerning the chemical reactions Zhan et al. (2023); Cortés (2017) stimulated by the electric field enhanced near plasmonic nanoparticles.
A photochemical reaction is initiated by the absorption of light and the creation of transient molecules in excited states whose properties differ greatly from the original molecules. These new chemical species can fall apart, change to new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules Fleming, Krueger, and Longworth (2023). So the optimization of the plasmonic photo-chemical reactions is important with the searching for conditions of the maximum field and absorption enhancement near metal nanoparticles Kelly et al. (2003). This paper contributes to such an optimization.
A large electric field enhancement is well-known near the sharp features of nanostructures, as "hot spots" Díaz-Núñez et al. (2019), or in a small gap between nanoparticles Kim et al. (2010). However, it may be difficult to make a sufficiently large number of "hot spots" or small gaps in nanostructures. So in this paper, we do not consider the field enhancement by "hot spots" or in gaps between nanoparticles. We restrict ourselves by the absorption enhancement in the near field of metal nanoparticles with the excited LPR.
We will pay attention on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and cause new physical effects. Spontaneous emission near metal nanoparticles has been investigated, for example, in Blanco and García de Abajo (2004). Collective effects as a super-radiance have been studied for the resonant absorbers, as atoms or molecules, in the near zone of the plasmonic particle Shahbazyan and Pustovit (2011); Protsenko and Uskov (2015); Protsenko et al. (2017). However, the absorption enhancement near the nanoparticle with an account for the spontaneous and collective emissions deserves more studies. The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, particularly, in the relation with recently studied plasmon-induced direct resonance energy transfer (PIRET) processes You, Ramakrishna, and Seideman (2018). A part of studies of the plasmon-enhanced absorption with the spontaneous and collective emissions to the LPR mode, with the direct energy transfer from the plasminic nanoparticle to the absorbing molecule, is done in the present paper.
Here we use the model for the resonant absorption near plasmonic nanoparticles similar to the laser model. The plasmonic nanoparticle makes a cavity for the electric field similar to the laser cavity, as in the plasmonic nanolaser Protsenko et al. (2005, 2008); Protsenko (2012). A part of our approach is similar to the analysis of the spontaneous emission and collective effects in lasers André et al. (2019); Protsenko and Uskov (2022).
The plasmonic mode provides good conditions for collective effects in the resonant absorption. Collective effects in the resonant medium are significant in the low-quality cavity André et al. (2019); Protsenko and Uskov (2022); Piccione, Felicetti, and Bellomo (2022). The LPR mode has a low quality, appropriate for collective effects. The particle surface is a good place for the resonant absorbers (atoms, molecules, q-dots, etc.).
In this paper, we show that the spontaneous emission into the LPR mode and collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation. We will find conditions for the best absorber excitation and bistability. These results are important for all branches of plasmonic chemistry, photovoltaic, and all applications related to the resonant excitation of absorbers near plasmonic nanoparticles.
We derive basic equations quantum-mechanically as Heisenberg operator equations. In the present paper, we solve these equations in the semi-classical approximation as it is common, for example, in the laser theory Sargent, Scully, and Lamb (1974). Fully quantum analysis of the equations will be done in the future.
We carry out calculations analytically. The analytical approach lets to find expressions illustrating the physical processes at the absorber excitation and the influence of the plasmonic particle on important parameters such as the polarization and the population of states of absorbers.
We consider the increase of the absorber excited state population as a criterion of absorption enhancement by the LPR.
In the next section, we describe the model, Hamiltonian, and equations of motion.
The expressions, characterizing the stationary absorption, as the absorption coefficients and the energy balance equation, are presented in section III. Several physical processes at the resonant absorption near the metal nanoparticle are explained with the help of equations found in section III.
Section IV shows conditions for the best plasmon-enhanced absorption and describes the bistability.
Examples of the absorption enhancement and the bistability are presented in section V at typical values of the nanoparticle and the absorber parameters.
Results are discussed in section VI, and the paper is finalized in the conclusion section VII.
II The model, Hamiltonian and equations of motion
II.1 The model
Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, being on the distance $r$ from the particle center on the dielectric shell as shown in Fig. 1. The absorber transition frequency is resonant with the dipole LPR mode of the nanoparticle.
Figure 1: Coordinate system, the nanoparticle 1 and the resonant absorber $i$ (a molecule or an atom) in the external electric field $\hat{\vec{E}}$. $r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the nanoparticle center to the absorber. $\left|00\right>_{i}$ is the absorber ground state. The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarization (or to the vector $\vec{e}_{i}$). The absorber state with the transition dipole momentum parallel to the axis $OO^{\prime}$ is not excited; such transition does not interact with the external and the nanoparticle fields. The insert shows the resonant absorbers randomly distributed on the nanoparticle shell.
Such resonance can be achieved for the nanoparticle of appropriate shape Kelly et al. (2003). For simplicity, we consider a spherical nanoparticle.
The particle and absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarized along the unit vector $\vec{e}$. The field $\hat{\vec{E}}$ is resonant with the LPR and the absorber transitions.
We consider the $i$-th absorber in the Cartesian coordinate system with the origin in the particle center. One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane made by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ directed from the nanoparticle center to the i-th absorber. The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third coordinate axis is along the line $OO^{\prime}$. Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of i-th absorber participate in the interaction with electric fields near the nanoparticle. Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) Fukui, Yonezawa, and Shingu (2004).
We will calculate the population of excited states of absorbers and examine how the interaction between the particle and absorbers affects the population.
We suppose a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell as shown on the insert in Fig. 1. In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near field leads to only the self-broadening of the absorber’s transitions Royer (1975). We include the self-broadening in the absorber transition width $\gamma$ and do not describe the absorber-absorber interaction explicitly.
II.2 Hamiltonian
Following Protsenko et al. (2005) we describe the nanoparticle as a dipole making harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
$\hat{\vec{p}}=p\vec{e}\left(\hat{b}{{e}^{-i{{\omega}_{0}}t}}+{{{\hat{b}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (1)
where $p$ is the dipole momentum matrix element, $\hat{b}$ is a Bose-operator describing collective (plasmonic) oscillations of the nanoparticle electrons in the electric field. Later we express $p$ through the nanoparticle polarizability Protsenko et al. (2005). For symmetry reasons, we suppose the same directions for the nanoparticle dipole and the external field polarization.
The dipole momentum operator of the $i$-th absorber resonant transitions is
${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right].$ (2)
Here $d$ is the dipole transition matrix element; absorber operators in Dirac notations ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
$H={{H}_{0}}-\left(\hat{\vec{p}}+\sum\limits_{i=1}^{N}{{{{\hat{\vec{d}}}}_{i}}}\right)\hat{\vec{E}}+{{\hat{V}}_{int}}+\hat{\Gamma}$ (3)
The first term in (3) describes free oscillations of dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field) and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
We suppose that the distance $r$ between absorbers and the particle center is much smaller than the LPR wavelength. So the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction. We take the dipole-dipole quasi-static interaction energy from Landau and Lifshitz (1982), replace classical dipoles with operators, and write
${{\hat{V}}_{int}}=\frac{1}{{{r}^{3}}}\sum\limits_{i=1}^{{{N}_{m}}}{\left[\hat{\vec{p}}{{{\hat{\vec{d}}}}_{i}}-3(\hat{\vec{p}}{{{\vec{e}}}_{ri}})({{{\hat{\vec{d}}}}_{i}}{{{\vec{e}}}_{ri}})\right]},$ (4)
where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle center to i-th absorber as shown in Fig. 1.
The external field is
$\hat{\vec{E}}={{E}_{0}}\vec{e}\left(\hat{a}{{e}^{-i{{\omega}_{0}t}}}+{{{\hat{a}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (5)
where $\hat{a}$ is the field amplitude Bose-operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ does not come to final results Lifshitz, Pitaevskii, and Berestetskii (1982)), $\eta$ is the refractive index of the nanoparticle environment Tkalya (2000). We suppose that distances between absorbers and the particle center are much smaller than the external field wavelength, so the phase (taken zero) of the external field is the same for all absorbers and the particle.
We substitute expressions (1), (2) into the sum in Eq (4), drop the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$ and represent
$\hat{V}_{int}=\frac{pd}{r^{3}}\sum\limits_{i=1}^{{{N}_{m}}}{\left\{{{{\hat{b}}}^{+}}\left[{{{\hat{\sigma}}}_{i}}-3\left({{{\hat{\sigma}}}_{i}}{{\cos}^{2}}{{\theta}_{i}}+{{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}\right)\right]+h.c.\right\}}$ (6)
Expression (6) can be simplified at a large number of absorbers $N_{m}\gg 1$. Summing over $i$ we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$. After such a simplification of Eq. (6) the Hamiltonian (3) written in the interaction picture Landau and Lifshitz (2013) and RWA approximation Sargent, Scully, and Lamb (1974) reads
$H=-{{\hat{a}}^{+}}\left(\hbar{{\Omega}_{p}}\hat{b}+\hbar{{\Omega}_{m}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}\right)-\hbar{{\Omega}_{int}}{{\hat{b}}^{+}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}+h.c.+\hat{\Gamma}$ (7)
where ${{\Omega}_{p}}={p{{E}_{0}}}/{\hbar}$, ${{\Omega}_{m}}={d{{E}_{0}}}/{\hbar}$ and ${{\Omega}_{int}}={pd}/2\hbar{{r}^{3}}$. The Hamiltonian (7) corresponds to the approximation when the polarization of absorber’s transitions is parallel to the external field polarization.
II.3 Equations of motion
We write Heisenberg equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from Hamiltonian (7). The equation for the plasmonic oscillation Bose-operator $\hat{b}$ is
$\dot{\hat{b}}=i{{\Omega}_{p}}\hat{a}+{{\Omega}_{int}}\hat{p}-\Gamma\hat{b}+(\sqrt{2\Gamma}){{\hat{b}}_{in}}$ (8)
where we introduce the polarization operator of all absorbers
$\hat{p}=i\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}.$ (9)
The last two terms are added to Eq. (8) following the input-output theory Courty and Reynaud (1992); Collett and Gardiner (1984), they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath. The stationary solution of Eq. (8) is
$\hat{b}=({i{{\Omega}_{p}}}/{\Gamma})\hat{a}+({{{\Omega}_{int}}}/{\Gamma})\hat{p}+(\sqrt{{2}/{\Gamma}}){{\hat{b}}_{in}}$ (10)
Using well-known properties of Dirac notations Lifshitz, Pitaevskii, and Berestetskii (1982) we come to commutation relations
$\left[{{{\hat{\sigma}}}_{i}},\hat{\sigma}_{i}^{+}\right]={{\hat{N}}_{0i}}-{{\hat{N}}_{1i}}$ (11)
where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are $i$-th absorber state population operators. Operators with different indexes $i$ commute. Using commutation relations (11), Hamiltonian (7) and polarisation (9) we write the equation of motion
$\dot{\hat{p}}=-\left({{\Omega}_{m}}\hat{a}+{{\Omega}_{int}}\hat{b}\right)\left({{{\hat{N}}}_{0}}-{{{\hat{N}}}_{1}}\right)-\gamma\hat{p}+\hat{F_{p}}$ (12)
where ${{\hat{N}}_{\alpha}}=\sum\limits_{i=1}^{N}{{{{\hat{N}}}_{\alpha i}}}$, $\alpha=0,1$ are operators of populations of states of all absorbers. The last two terms in Eq. (12) describe the absorber polarization decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarization baths.
Taking commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7) and summing over all absorbers we obtain the equation of motion
${{\dot{\hat{N}}}_{0}}=\left({{\Omega}_{m}}{{{\hat{a}}}^{+}}+{{\Omega}_{int}}{{{\hat{b}}}^{+}}\right)\hat{p}+h.c.+{{\hat{N}}_{1}}/\tau+{{\hat{F}}_{0}}$ (13)
where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population baths. The population operator of the upper states of absorbers is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
III The stationary absorption
From here we use the semi-classical approach as, for example, in the laser theory Sargent, Scully, and Lamb (1974). So we neglect fluctuations, drop the bath operators, and replace operators with c-numbers (removing hats from notations). Calculating the semi-classical stationary solution of Eqs. (10) and (12) we find
$p=-\frac{\left({{{\Omega}_{m}}}/{\gamma}+{i{{\Omega}_{int}}{{\Omega}_{p}}}/{\Gamma\gamma}\right)a\left({{N}_{0}}-{{N}_{1}}\right)}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}},$ (14)
where ${{N}_{int}}=\gamma\,\Gamma/\Omega_{int}^{2}$.
Eq. (14) helps us to understand physical processes in the absorber-particle interaction. The stationary polarization of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$. The particle near absorbers interacts with the external field with the rate $\Omega_{p}$, accepts the polarization, re-emits the field, and, by the re-emitted field, "transfers" the polarization (and the energy) to absorbers with the rate $\Omega_{int}$. Such polarization and the energy transfer is the plasmon-induced resonance energy transfer (PIRET) You, Ramakrishna, and Seideman (2018). PIRET is suppressed by the polarization decay: in the particle with the rate $\Gamma$ and in absorbers with the rate $\gamma$. This is why the PIRET particle-absorber polarization transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq. (14). The multiplier $i$ in the last expression and in Eq. (14) is due to the well-known $\pi/2$ phase shift at the resonant interaction of oscillators Lifshitz and Landau (1982).
The absorber-particle interaction is a reciprocal process: the particle transfers the polarization to absorbers and vice-versa: absorbers transfer the polarization to the particle through the near field, causing the absorber polarization reduction. Such a reduction is described by the factor $1+\left({{N}_{0}}-{{N}_{1}}\right)/N_{int}$ in the denominator of Eq. (14).
In order to understand the physics and to characterize the stationary absorption we write, using Eq. (13), the energy conservation law: the rate $N_{1}/\tau$ of the energy relaxation from the upper states of absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle and absorbers through plasmonic mode $\sim\Omega_{int}$
$\frac{{{N}_{1}}}{\tau}=\left(-{{\Omega}_{m}}+\frac{{{i{\Omega}_{int}\Omega}_{p}}}{\Gamma}\right){{a}^{*}}p+c.c.-\frac{2\Omega_{int}^{2}}{\Gamma}\left\langle{{{\hat{p}}}^{+}}\hat{p}\right\rangle,$ (15)
[…]
$\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}=\hat{N}_{1i}$ (16)
and assuming a large number of absorbers $N_{m}\gg 1$ we write
$\left\langle{{{\hat{p}}}^{+}}\hat{p}\right\rangle={{N}_{1}}+\sum\limits_{i\neq j}{\left\langle\hat{\sigma}_{i}^{+}{{{\hat{\sigma}}}_{j}}\right\rangle}\approx{{N}_{1}}+{{\left|p\right|}^{2}}$ (17)
The term $N_{1}$ in Eq. (17) describes the spontaneous emission, and the second term - a collective emission of absorbers to the plasmonic mode. These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory André et al. (2019); Protsenko and Uskov (2022). We find from Eqs. (14), (15) and (17)
$\frac{{{N}_{1}}}{\tau}=$ ${\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-}-$
$\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right).$
where $n=|a|^{2}$ is the mean photon number in the volume $V$ of the external field. We express $n/V$ through the external field intensity later.
We introduce the absorption coefficients: ${{g}_{m}}={2\Omega_{m}^{2}}/{\gamma}$ of an absorber, ${{g}_{p}}={2\Omega_{m}^{2}}/{\Gamma}$ of the particle and the spontaneous emission rate of an absorber to the plasmonic mode $\tilde{\gamma}_{sp}=2\Omega_{int}^{2}/\Gamma$. Using Eq. (14) we find
$\frac{2\Omega_{int}^{2}}{\Gamma}{{\left|p\right|}^{2}}=n\frac{\left({{g}_{m}}+{{g}_{p}}/{{N}_{int}}\right){{{\tilde{\gamma}}}_{sp}}}{2\gamma}{{\left(\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right)}^{2}}$ (19)
and represent Eq. (III) in a compact form
$\frac{{{N}_{1}}}{\tau}=n{{\tilde{g}}_{m}}({{N}_{1}})\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-{{\tilde{\gamma}}_{sp}}{{N}_{1}}$ (20)
where the nonlinear absorption coefficient
${{\tilde{g}}_{m}}({{N}_{1}})=\left({{g}_{m}}+\frac{{{g}_{p}}}{{{N}_{int}}}\right)\left[1-\frac{{{{\tilde{\gamma}}}_{sp}}}{2\gamma}\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right].$ (21)
With the help of Eqs. (19) – (21) we identify processes affecting the absorber excited states population. The first process is the resonant absorption of the external field with the absorption coefficient $g_{m}$. The second process is the resonant absorption increase due to the nanoparticle near field. An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to absorbers by the near field with the transfer coefficient $1/N_{int}$. So the linear (i.e. independent on $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$. The nonlinear term $\sim\tilde{\gamma}_{sp}$ in Eq. (21) is due to the collective emission of absorbers to plasmonic mode. The reason for the suppression factor $1+(N_{0}-N_{1})/N_{int}$ in the denominator of Eq. (20) is the transfer of the polarization and the energy from absorbers to the particle. The upper states population decays with the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq. (20).
IV Plasmon-enhanced absorption conditions and bistability
The resonant absorption in plasmonic mode is a nonlinear process. Besides the well-known resonant transition saturation, the nonlinearity is caused by the collective interaction of absorbers and the particle through the plasmonic mode leading to the energy exchange between absorbers and the particle. We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability in absorption and affect conditions for the best excitation of absorbers. Let us introduce some parameters.
IV.1 Parameters
We express
${{g}_{p}}n=\frac{2\Omega_{p}^{2}n}{\Gamma}=\frac{2n}{\Gamma}{{\left(\frac{p{{E}_{0}}}{\hbar}\right)}^{2}}=\frac{2n}{\Gamma}{{\left(\frac{p}{\hbar}\right)}^{2}}\frac{1}{{{\eta}^{2}}}\frac{2\pi\hbar{{\omega}_{0}}}{V}=\frac{4\pi n\omega_{0}\alpha_{p}}{\eta^{2}V}.$
Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nano-particle Protsenko et al. (2005). The external field intensity is $I=({c\eta}/{4\pi})\left\langle{{{\hat{\vec{E}}}}^{+}}\hat{\vec{E}}\right\rangle={c\hbar{{\omega}_{0}}n}/{V\eta}$. Here and below $c$ is the speed of light in vacuum. We take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second. So the number of photons in the external field per the unit of volume is ${n}/{V}={\eta P}/{c}$ and
${{g}_{p}}n={{\alpha}_{p}}\frac{4\pi{{\omega}_{0}}}{\eta}P=\frac{8{{\pi}^{2}}{{\alpha}_{p}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{p}}P,$ (22)
where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the absorbing field wavelength in vacuum, $\sigma_{p}$ is the particle resonant absorption cross-section. Similar way we find
${{g}_{m}}n=\frac{8{{\pi}^{2}}{{\alpha}_{m}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{m}}P,$ (23)
where $\alpha_{m}$ is the absolute value of the polarizability and $\sigma_{m}$ is the absorption cross-section of the two-level resonant absorber. It is convenient to represent the resonant absorber’s polarizability
$\alpha_{m}=3\gamma_{sp}/4\eta k_{0}^{3}\gamma,$ (24)
in terms of natural linewidth $\gamma_{sp}$ of the resonant transition Scully and Zubairy (1997); $k_{0}=\omega_{0}/c$ is the absorbed field wave-vector in vacuum. Using the parameters introduced above we express
$\frac{1}{{{N}_{{int}}}}\equiv\frac{\Omega_{{int}}^{2}}{\gamma\Gamma}=\frac{{{p}^{2}}}{\hbar\Gamma}\frac{{{d}^{2}}}{\hbar\gamma}\frac{1}{4{{r}^{6}}}=\frac{{{\alpha}_{p}}{{\alpha}_{m}}}{4{{r}^{6}}},$
$\tilde{\gamma}_{sp}=\frac{2\Omega_{int}^{2}}{\Gamma}=\frac{\alpha_{p}\alpha_{m}}{2r^{6}}\gamma=\frac{2\gamma}{N_{int}}.$
IV.2 Conditions for the absorption enhancement
We derive from the energy conservation law (20) and (21)
${{N}_{1}}={{\sigma}_{eff}}({{N}_{1}})P\tau\left({{N}_{m}}-2{{N}_{1}}\right),$ (25)
where ${{\sigma}_{eff}}({{N}_{1}})$ is a nonlinear (i.e. depended on $N_{1}$) cross-section of absorption
$\sigma_{eff}(N_{1})=\frac{\xi_{\sigma}\sigma_{m}}{[1+(N_{m}-2N_{1})(\alpha_{p}\alpha_{m}/4r^{6})]^{2}}.$ (26)
In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross-section enhancement factor
$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}$ (27)
The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to absorbing molecules. Such transfer is described by the factor $\alpha_{p}^{2}/4r^{6}$. The denominator of Eq. (27) is greater than 1 because of the spontaneous emission to the LPR mode. Such spontaneous emission reduces the absorption cross-section. $\sigma_{eff}$ is reduced relatively to $\xi_{\sigma}\sigma_{m}$ because of the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from emitters to the particle. Both these processes make the denominator in Eq. (26) greater than 1.
Conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27). The absorption enhancement is, when $\sigma_{eff}(N_{1})>\sigma_{m}$. Suppose, in the beginning, that the denominator in Eq. (26) is close to 1, so the collective emission to the LPR mode and the energy transfer from absorbers to the nanoparticle are relatively small. This is true at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number $N_{m}$ of absorbers, when
$N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\leq 1.$ (28)
When Eq. (28) is true, the absorption cross-section is enhanced at $\xi_{\sigma}>1$ which is when
$\alpha_{p}/\alpha_{m}>2\gamma\tau,\hskip 14.22636pt$ (29)
or, if we insert the expression (24) for $\alpha_{m}$ to (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$. The relation (29) is a necessary condition for the plasmon-stimulated absorption enhancement (found in the semiclassical approach). We note that $2\gamma\tau\geq 1$, so it must be $\alpha_{p}>\alpha_{m}$ for the absorption enhancement. So absorption enhancement is possible when the nanoparticle absorbs the field better than an emitter, which has been mentioned, for example, in Cortés (2017).
However the nanoparticle absorption can not be too large. Otherwise, the absorption by molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from molecules to the particle. So combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the the absorption enhancement by LPR is possible
$2\gamma\tau\alpha_{m}\equiv\alpha_{p}^{min}\leq\alpha_{p}\leq\alpha_{p}^{max}\equiv 2r^{6}/N_{m}\alpha_{m}.$ (30)
[…]
It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as far as absorbers are in the nanoparticle near the field zone). However, the magnitude of the enhancement depends on $r$, as we will see.
IV.3 Bistability in the stationary absorption
Suppose the number $N_{m}$ of absorbing molecules is so large that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$. Then the absorption is strongly nonlinear, because of the collective effects, the energy transfer between absorbers and the particle, which makes a significant effect of $N_{1}$-dependent denominator in Eq. (26). The denominator is decreased with $N_{1}$, so the absorption is increased with $N_{1}$ at certain parameter values. As a result the low and the high absorption regimes are possible, they overlap at certain parameters making the bistability in absorption.
In order to investigate the bistability we introduce dimensionless absober-nanoparticle interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$. With these parameters we write a cubic equation on $n_{1}$
${{{n}_{1}}}/({1-2{{n}_{1}}})+2{{n}_{1}}A+{{n}_{1}}\left(1-2{{n}_{1}}\right){{A}^{2}}={{P}_{eff}},$ (31)
instead of Eq. (25). Proceeding a standard analysis of the catastrophe theory Poston and Stewart (1996) we find that Eq. (31) has three solutions if the number $N_{m}$ of absorbers is so large that
$A>A_{cr}=8.$ (32)
Note that the condition (32) is opposite to the small collective emission and the small absorber to particle energy transfer condition (28). Thus the collective emission and the absorber to particle energy transfer must be sufficiently large for the bistability.
The bistability domain is restricted by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
$P_{\pm}(A)={{{n}_{\pm}}}/({1-2{{n}_{\pm}}})+2{{n}_{\pm}}A+{{n}_{\pm}}(1-2{{n}_{\pm}}){{A}^{2}}$ (33)
where ${{n}_{\pm}}(A)=\left(3\pm\sqrt{1-8/A}\right)/8$. The bistability domain (33) is shown in Fig. 2.
Figure 2: The absorption bistability is for values of $A$ and $P_{eff}$ between the curves $P_{\pm}(A)$.
V Examples of absorption enhancement and the bistability
In examples we consider a spherical silver nanoparticle with a polymer shell in the water. The refractive index of most polymers is close to the glass refractive index $1.55$ Zhang et al. (2020). The water refractive index is $1.33$. For simplicity, we take the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
Nanoparticle resonant polarizability is calculated in Appendix taking into account the radiative losses Bottcher (1973); Protsenko and Uskov (2012) and the losses due to collisions of the metal electrons with the nanoparticle surface Kik and Brongersma (2007); Protsenko and Uskov (2012). We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half-wigth $\Gamma\approx 80$ THz. We take the nanoparticle radius $r_{p}=13$ nm corresponding to the maximum LPR quality factor $Q=29$. Resonant absorbers are on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm. We take the minimum shell thickness $5$ nm because of a rapid de-population of the absorber excited states (fluorescence quenching) at $\Delta r<5$ nm due to the absorber interaction with the high-order (quadrupole etc.) plasmonic modes Dulkeith et al. (2005). The number of absorbers on the shell is large $N_{m}\geq 10$.
We consider the stationary population $n_{1}$ of the upper level of the absorber resonant transition as a characteristic of the absorption. We compare $n_{1}$ with and without the nano-particle, so we examine the LPR-stimulated absorption enhancement (or suppression) at various numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper-level de-population time $\tau$ and the absorber transition half-width $\gamma$.
[…]
Figure 3: (a) The stationary upper-level population per absorber versus the normalized external field intensity at various number of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3). Curve 4 is for the absorber upper-level population in free space. (b) The ratio $R$ of the emitter upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a. The unstable stationary absorption corresponds to dashed parts of curves 3.
Curves 3 in Figs. 3a, b are for $N_{m}=10$, when $\alpha_{p}/\alpha_{p}^{max}\approx 1.8\sim 1$ and both conditions (28) and (29) (or (30)) are satisfied. We see from curves 1 on Figs. 3 that the LPR absorption is enhanced, relative to the free space absorption, at any input field intensity. The maximum enhancement is about 2.5 –3 times, respectively to free space, at a small field intensity far from the absorber transition saturation.
We increase the number of absorbers. Curves 2 in Figs. 3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and the absorber-to-particle energy transfer is not satisfied. We see the absorption suppression by LPR at small $\sigma_{m}P\tau<1.6$. Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the absorber transition saturation. Note that $A=0.97<A_{cr}$ for curves 2, so there is no absorption bistability.
We increase $N_{m}$ more and take $N_{m}=120$ for curves 3 in Figs. 3, so $A=11.6>A_{cr}=8$, the condition (32) is satisfied and we see the bistability in curves 3. The bistability is between regimes with the low and the high absorption. The LPR suppresses the absorption at a small and enhances it at a large external field intensity. The unstable stationary solution corresponds to the dashed parts of curves 3.
Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=120$ and different thickness of the nanoparticle shell $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3). The rest of the parameters is the same as for Figure 3. The absorber without the nanoparticle corresponds to curve 4. The dashed parts of curves mean the unstable stationary solution. The bistability disappears with the increase of the shell thickness $\Delta r$.
Figure 4 shows $n_{1}$ and $R$ dependencies on the normalized external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different thicknesses of the nanoparticle shell $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3). We see that the bistability disappears with only 1 nm increase of the nanoparticle shell thickness. For each shell thickness, the absorption is suppressed at small input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
Let us see what happens with the decrease of the absorber state de-population rate (i.e. the increase $\tau$).
Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=10$ and different absorber upper level de-population time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3). Curve 4 is for an absorber without the nanoparticle. The shell thickness $\Delta r=5$ nm, the rest of parameters is the same as for Figure 3. The upper-level population and the absorption are reduced with the increase of $\tau$.
The absorption enhancement by LPR is reduced with the increase of $\tau$, from curves 1 to curves 3 in Fig 5, because of the growing of the spontaneous emission to the LPR contribution to the population decay.
Suppose we have a large absorber upper state de-population time $\tau=1/2\gamma_{sp}=0.5$ ns; such a de-population is, for example, because of the spontaneous emission to free space with the typical optical domain rate $\gamma_{sp}=1$ GHz Scully and Zubairy (1997). We assume the absorber transition half-width $\gamma=100$ GHz. Then $\alpha_{p}/\alpha_{m}=20<2\gamma\tau=100$ so the condition (29) is not satisfied and there is no absorption enhancement by the LPR at any external field intensity. Meanwhile, the bistability in absorption exists, as shown by curve 3 in Figure 6, when the bistability condition (32) is true.
Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the suppression of absorption by the spontaneous emission to LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz. The number of absorbers $N_{m}=10$ is for curve 1; $N_{m}=16$ is for curve 2 and $N_{m}=30$ is for curve 3; curve 4 is for the absorber in free space. The bistability condition (32) is satisfied and the bistability in absorption exists for curve 3. The dashed part of curve 3 corresponds to the unstable stationary state.
VI Discussion
An electric field enhancement near plasmonic nanoparticles increases the resonant absorption of atoms, molecules, semiconductor media, etc. The estimation (27) shows the growth of the absorption cross-section near plasmonic particle in a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq. (27)). Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it, at some conditions, smaller than in the free space. In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about the factor 3 (see curve 1 in Figure 3b), with the spontaneous emission to the LPR mode taken into account.
The maximum absorption enhancement is found for a small external field when the absorbing transition is not saturated. A noticeable absorption enhancement is possible at a fast (on ps time scale) de-population of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is about or less than the de-population rate due to other processes as, for example, the decay of absorber excitation to the non-resonant high-order plasmonic modes Dulkeith et al. (2005). When the population decay is not so fast, the relative suppression of the absorber excitation by the radiation to the LPR is large and we do not have the absorption enhancement, as shown in Figure 6. The condition (29) must be satisfied for the LPR absorption enhancement. Spontaneous and collective emissions to the LPR mode are important factors, they can not be neglected in the resonant absorption near metal nanoparticles.
The spontaneous emission enhancement by the cavity mode is well-known in the laser theory Oraevskii (1994) as a Pursell effect Purcell (1946). The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity enhances the spontaneous emission similar to a laser cavity.
When a few resonant absorbers are placed near metal nanoparticles, a collective emission from absorbers to the plasmonic mode is possible. A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and emitted again by the collective emission. This is a Dicke effect Gross and Haroche (1982) with the probability (and, therefore, the rate) larger than for the photon emission by a single atom. The collective emission de-populates the absorber’s upper state. Collective emission to the LPR mode is similar to the collective emission to a lasing mode Leymann et al. (2015); Jahnke et al. (2016); André et al. (2019); Protsenko and Uskov (2022). The collective emission is a nonlinear process, its rate depends on the emitting state populations. In particular, the collective emission is suppressed at the saturation of the emitting transition.
The particle-absorber near-field interaction is a mutual process: the metal nano-particle LPR mode excites the polarization and transfers the energy to the absorber through the near field and vice-versa: absorbers excite the particle polarisation through the near field. Such an energy transfer is the origin of the suppression factor in the denominator in Eq. (14) for the absorber polarisation. The collective emission and the emitter-particle interaction depend on the absorber states population. Such a dependence leads to a positive feedback: the absorption is increased with the upper-level population at certain conditions. The feedback is fast: note the square in the nonlinear denominator in Eq. (26). So, at certain parameter values, the feedback nonlinearity "overtakes" the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity. When $n_{1}$ decreases, the stationary absorption is unstable, see the dashed parts of curves in Figs. 3, 4, and 6. Such anomalous absorption causes absorption bistability. The bistability exists at the condition (32) which can be satisfied at the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6. With the LPR absorption enhancement, the bistability is between the absorption suppression and the absorption enhancement regimes, see Figures 3 and 4.
The optical bistability is possible in plasmonic nanolaser with external signal Protsenko et al. (2008). The model of Protsenko et al. (2008) involves a single emitter near the plasmonic nanoparticle and does not include the spontaneous and collective emissions to the LPR mode.
The physical nature of a well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) Koch, Schmidt, and Haug (1985) is the resonant transition saturation. This is different from the physics of the bistability in the resonant absorption in the LPR near metal nanoparticle described here. The absorptive bistability Koch, Schmidt, and Haug (1985) is not related to the collective emission and the near-field interaction of absorbers with the metal nanoparticle.
The nanoparticle with resonant emitters is, probably, one of the smallest systems with optical bistability with the size much smaller than the optical wavelength.
Here we carry out the semi-classical analysis of the resonant absorption near a metal nanoparticle. Only the term of the spontaneous emission into the LPR mode in Eq. (17) is introduced according to quantum mechanical relation (16). We note that the term of the spontaneous emission to the lasing mode appears in the laser theory the same way P. R. Rice  and Carmichael (1994); André et al. (2019). More detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber states populations. Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode Maier (2006), which keeps only one or a few photons. Such an analysis will be carried out in the future.
VII Conclusion
We investigate the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) near plasmonic nanoparticles, taking into account the spontaneous and collective emission from absorbers to the localized plasmon resonance (LPR) mode of the particle. We derive quantum equations for the particle and the resonant absorber polarisation and the absorbing state populations. We analyze the structure of equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
We find that the absorption enhancement by the LPR is possible under certain conditions. The absorption suppression is possible as well due to the spontaneous, collective emissions of absorbers to the LPR mode and the energy transfer from absorbers to the particle. We find conditions necessary for the absorption enhancement. One condition is a fast, of the pico-second time scale, de-population of the excited absorber states, for example, by some non-radiative processes.
We find conditions and investigate the bistability in the resonant absorption near the metal nanoparticle. The bistability is caused by the collective emission and the nonlinear interaction between the particle and absorbers through the LPR near field. The bistability appears at a large number of absorbers near the nanoparticle and possible with the absorption enhancement or suppression.
Results can be used everywhere, when the resonant absorption near metal nanoparticles is used, for example, for the plasmon-stimulated photo-catalysis Zhang et al. (2013) or in the plasmonic photo-voltaic Jang et al. (2016).
Acknowledgements.
I.E.P., A.V.U. and N.V.N. acknowledge the Russian Science Foundation (Grant No.20-19-00559) for the support.
Appendix. Nanoparticle resonant polarisability
Polarizability $\alpha_{p}$ of the spherical nanoparticle of the radius $r_{p}$ in the electric field with the wavelength $\lambda$, in the dielectric environment with the refractive index $\eta$ is Protsenko and Uskov (2012)
$\alpha_{p}(\lambda,r_{p})=r_{p}^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{[1-iR_{rad}(\lambda,r_{p})][\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2]}$ (34)
where $\varepsilon_{p}$ is the dielectric function of the material of the nanoparticle and the factor
$R_{rad}(\lambda,r_{p})=\frac{16\pi^{3}}{3}\left(\frac{\eta r_{p}}{\lambda}\right)^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2}$ (35)
describe the radiation losses.
The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ takes into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary Protsenko and Uskov (2012); Uskov et al. (2014)
$\varepsilon_{p}(\lambda,r_{p})=\varepsilon_{0}-\frac{(\lambda/\lambda_{p})^{2}}{(1+i\lambda/\lambda_{c})(r_{c}/r_{p}+1)}.$ (36)
[…]